# The NEID Earth Twin Survey. III. Survey Performance After Three Years on Sky

### v1

@@ 5.3.5 HD 55575
For the star HD 55575, our initial test prefers a model without a trend. In both searches, we identify a 52.5-day signal with m s-1 and . Rosenthal et al. (2021) report the detection of a RV signal at this period which they attribute to the stellar rotation period. However, there is no published rotation period for this star, and neither the HIRES S-index measurements presented by Rosenthal et al. (2021) nor any of the NEID activity indicators presented herein exhibit any periodicity at 52 days or at any aliases thereof. A more thorough investigation is warranted to re-assess the classification of this signal.
@@ 5.3.7 HD 146233
Using RV measurements of HD 146233 from HARPS-N, HIRES, PFS, and UCLES in a blind search analysis similar to the one implemented here, Laliotis et al. (2023) identified a candidate planetary signal with days, m s-1, and . For our search that allows for a trend, we detect this same 19.9-day signal ( m s-1, ) alongside a long-term linear trend of m s-1 yr-1. The search results are shown in Figure 11. The sign and slope of the long-term trend are consistent with the known long-term activity cycle for HD 146233 as well as a moderate correlation between the NEID RVs and SHK measurements (Spearman’s (Run 1) and (Run 2)). In our search without a trend, we do observe a peak in the BIC periodogram at 19.9 days, but this peak does not exceed the false alarm threshold.
@@ 5.3.9 HD 126053
HD 126053 is a Solar-type star with a 26-day rotation period (Hempelmann et al., 2016). Our initial test prefers a model with no trend; in both searches, we detect a 10.9-day RV signal with m s-1 and . While we do see power near the stellar rotation period in multiple activity indicators, including SHK and H, the rotation signal is absent from the RVs.
@@ 5.3.12 HD 185144
HD 185144 ( Dra) is an early K-dwarf with no known planetary companions, though the star does have a well-characterized, -yr stellar activity cycle (Baum et al., 2022; Isaacson et al., 2024). The search with a trend adopts a quadratic model with linear and quadratic trend terms of m s-1 yr-1 and m s-1 yr-2, and no signals are detected. Without a trend, the search returns a significant signal at d, with orbital eccentricity and semi-amplitude m s-1. Significant long-period power is present in the stellar activity indicators, corresponding to a change in the overall activity level as the star enters an activity minimum. The quadratic trend and the detected eccentric signal could be explained by the RV response to the stellar activity cycle, but further investigation is warranted.
@@ 5.3.13 HD 68017
HD 68017 is a G-dwarf with an M-dwarf companion on a very close orbit (0.59”, projected separation 13 au; Crepp et al., 2012). In both search cases ("blind" with best-fit linear and quadratic trend terms of m s-1 yr-1 and m s-1 yr-2, and "informed" of the known companion parameters) a 118-day signal is identified with m s-1, and m s-1, for the blind and informed saerches, respectively. We do not see peaks at this period for any of the activity indicators, nor are the S-index and RVs linearly correlated. In addition, a signal at this period is also present in the RV residuals to the Giovinazzi et al. (2025) joint RV+astrometric stellar orbit fit, although this signal does not exceed the false alarm threshold. This signal merits further investigation to determine whether it is of planetary origin.
@@ 5.4.1 HD 26965
Multiple studies have investigated the nature of a 42-day signal appearing in the RVs of HD 26965. While this was initially reported as a candidate super-Earth, HD 26965 b by Díaz et al. (2018) and subsequently confirmed by Ma et al. (2018), other studies have posited that the RVs are driven by stellar activity (Zhao et al., 2022; Laliotis et al., 2023). Most recently, multiple independent analyses of spectrum-level variations in the Run 1 NEID data have firmly classified this as a rotationally-modulated activity signal (Burrows et al., 2024; Siegel et al., 2024; Gilbertson et al., 2024).
@@ 5.4.2 HD 38858
Mayor et al. (2011) reported the detection of a super-Earth around HD 38858 on an eccentric () orbit with a period of 407 days, but subsequent analyses show that this was an alias of the stellar activity cycle (Kennedy et al., 2015; Flores et al., 2018).
@@ 6.1 Sensitivity to known exoplanets
Our analysis in Sections 4 and 5 demonstrates that with just 3 years of dedicated NEID observations, our survey is approaching the same sensitivity as all cumulative RV monitoring of these same target stars over the past few decades. We successfully detect 10 of the 11 short-period () planets (period less than half the NEID baseline) in our sample (Figure 12), even in cases where no prior information is provided about long-period signals. In addition, we note that all of these detections, including several with m s-1, were achieved without any careful treatment (i.e., modeling, mitigation) of stellar variability. This is a testament to the advantages of selecting quiet, inactive target stars for the survey and observing with a highly stabilized instrument.
None of the five long-period planets (period greater than half the NEID baseline) are recovered (Figure 12), but this is unsurprising given that in most cases NETS has only covered a small fraction of their orbits. This will continue to be true through the end of the five-year survey, where only HD 186427 b ( d) will end up with full coverage over multiple orbits. Still, as Giovinazzi et al. (2025) show, NEID data can be combined with archival data over longer baselines to place stringent constraints on the orbits of long-period planets. Further, Figure 12 illustrates that the survey sensitivity is presently limited by the observing baseline, not by the RV precision or the intrinsic variability of the target stars. As the survey progresses, we expect to be able to robustly detect planets with m s-1 out to periods of several hundred days.
@@ 6.2 Targets on ExEP List
The 2020 Decadal Survey of Astronomy and Astrophysics (National Academies of Sciences, Engineering, and Medicine, 2021) recommended the development and construction of the Habitable Worlds Observatory (HWO), a flagship space telescope with high contrast direct imaging and spectroscopy capabilities from UV to IR wavelengths. This telescope will carry out a survey of nearby stars to search for and and characterize the atmospheres of at least 25 habitable-zone, terrestrial exoplanets. HWO is presently expected to be able to accomplish this through an uninformed search. However, information gleaned from EPRV surveys such as NETS has the potential to greatly increase the efficiency and scientific return of the mission. As Morgan et al. (2021) show, partial prior knowledge of Earth-like planet orbits from RV measurements can be used to improve HWO survey yield (i.e., the number of characterized planets) by up to 50% and accelerate the detection of key atmospheric biomarkers. Further, recent work by Kane et al. (2024) has highlighted the obstacles posed by as-yet undetected massive planets orbiting in or near the habitable zone. Planets as small as can dynamically preclude stable orbits for terrestrial planets through much of the habitable zones of their host stars, rendering these targets unfit for the HWO survey. But for the majority of high priority HWO target stars identified by Mamajek & Stapelfeldt (2024), existing RV data are insufficient to identify or rule out such perturbers (Laliotis et al., 2023; Harada et al., 2024).
@@
Independent of these challenges, accurate interpretation of atmospheric spectra collected by HWO or similar missions will require precise knowledge of the exoplanets’ physical properties. In particular, insufficient constraints on the planetary mass can lead to large uncertainties in the inferred atmospheric parameters (Batalha et al., 2019) and incorrect assessments of habitability and of the presence or absence of biosignatures (Damiano et al., 2025). Damiano et al. (2025) recommend a mass precision of better than 10% to guarantee reliable results from reflected light spectroscopy of Earth-like planets.
@@
The NETS target list includes 23 high-priority HWO target stars, including many with known planets (HD 115617, HD 143761, HD 190360, HD 86728, HD 95735) and others with promising, low-amplitude candidate RV signals identified in this work and previous works (HD 4628, HD 10780, HD 146233). It is essential that these stars continue to be monitored with NETS and other EPRV surveys to guarantee that we have an accurate census of their planetary systems and that these systems are well characterized in preparation for HWO. At the very least, establishing and maintaining an EPRV baseline now will greatly accelerate mass measurements for Earth-like planets independently detected by HWO. NETS data can also serve an important role in continuing to study the impact of stellar variability on the spectra of these host stars; while the NETS targets are quiet, they are not silent, and activity-driven spectral variations are well known to complicate exoplanet spectroscopy measurements (Rackham et al., 2018, 2019).

### v2

@@ 5.3.5 HD 55575
For the star HD 55575, our initial test prefers a model without a trend. In both searches, we identify a 52.5-day signal with m s-1 and . L. J. Rosenthal et al. (2021) report the detection of a RV signal at this period which they attribute to the stellar rotation period. However, there is no published rotation period for this star, and neither the HIRES S-index measurements presented by L. J. Rosenthal et al. (2021) nor any of the NEID activity indicators presented herein exhibit any periodicity at 52 days or at any aliases thereof. A more thorough investigation is warranted to re-assess the classification of this signal.
@@ 5.3.7 HD 146233
Using RV measurements of HD 146233 from HARPS-N, HIRES, PFS, and UCLES in a blind search analysis similar to the one implemented here, K. Laliotis et al. (2023) identified a candidate planetary signal with days, m s-1, and . For our search that allows for a trend, we detect this same 19.9-day signal ( m s-1, ) alongside a long-term linear trend of m s-1 yr-1. The search results are shown in Figure 11. The sign and slope of the long-term trend are consistent with the known long-term activity cycle for HD 146233 as well as a moderate correlation between the NEID RVs and SHK measurements (Spearman’s (Run 1) and (Run 2)). In our search without a trend, we do observe a peak in the BIC periodogram at 19.9 days, but this peak does not exceed the false alarm threshold.
@@ 5.3.9 HD 126053
HD 126053 is a Solar-type star with a 26-day rotation period (A. Hempelmann et al., 2016). Our initial test prefers a model with no trend; in both searches, we detect a 10.9-day RV signal with m s-1 and . While we do see power near the stellar rotation period in multiple activity indicators, including SHK and H, the rotation signal is absent from the RVs.
@@ 5.3.12 HD 185144
HD 185144 ( Dra) is an early K-dwarf with no known planetary companions, though the star does have a well-characterized, -yr stellar activity cycle (A. C. Baum et al., 2022; H. Isaacson et al., 2024). The search with a trend adopts a quadratic model with linear and quadratic trend terms of m s-1 yr-1 and m s-1 yr-2, and no signals are detected. Without a trend, the search returns a significant signal at d, with orbital eccentricity and semi-amplitude m s-1. Significant long-period power is present in the stellar activity indicators, corresponding to a change in the overall activity level as the star enters an activity minimum. The quadratic trend and the detected eccentric signal could be explained by the RV response to the stellar activity cycle, but further investigation is warranted.
@@ 5.3.13 HD 68017
HD 68017 is a G-dwarf with an M-dwarf companion on a very close orbit (0.59”, projected separation 13 au; J. R. Crepp et al., 2012). In both search cases ("blind" with best-fit linear and quadratic trend terms of m s-1 yr-1 and m s-1 yr-2, and "informed" of the known companion parameters) a 118-day signal is identified with m s-1, and m s-1, for the blind and informed saerches, respectively. We do not see peaks at this period for any of the activity indicators, nor are the S-index and RVs linearly correlated. In addition, a signal at this period is also present in the RV residuals to the M. R. Giovinazzi et al. (2025) joint RV+astrometric stellar orbit fit, although this signal does not exceed the false alarm threshold. This signal merits further investigation to determine whether it is of planetary origin.
@@ 5.4.1 HD 26965
Multiple studies have investigated the nature of a 42-day signal appearing in the RVs of HD 26965. While this was initially reported as a candidate super-Earth, HD 26965 b by M. R. D´ıaz et al. (2018) and subsequently confirmed by B. Ma et al. (2018), other studies have posited that the RVs are driven by stellar activity (L. L. Zhao et al., 2022; K. Laliotis et al., 2023). Most recently, multiple independent analyses of spectrum-level variations in the Run 1 NEID data have firmly classified this as a rotationally-modulated activity signal (A. Burrows et al., 2024; J. C. Siegel et al., 2024; C. Gilbertson et al., 2024).
@@ 5.4.2 HD 38858
M. Mayor et al. (2011) reported the detection of a super-Earth around HD 38858 on an eccentric () orbit with a period of 407 days, but subsequent analyses show that this was an alias of the stellar activity cycle (G. M. Kennedy et al., 2015; M. Flores et al., 2018).
@@ 6.1 Sensitivity to known exoplanets
Our analysis in Sections 4 and 5 demonstrates that with just 3 years of dedicated NEID observations, our survey is approaching the same sensitivity as all cumulative RV monitoring of these same target stars over the past few decades. We successfully detect 10 of the 12 short-period () planets (period less than half the NEID baseline) in our sample (Figure 12), even in cases where no prior information is provided about long-period signals. In addition, we note that all of these detections, including several with m s-1, were achieved without any careful treatment (i.e., modeling, mitigation) of stellar variability. This is a testament to the advantages of selecting quiet, inactive target stars for the survey and observing with a highly stabilized instrument.
None of the five long-period planets (period greater than half the NEID baseline) are recovered (Figure 12), but this is unsurprising given that in most cases NETS has only covered a small fraction of their orbits. This will continue to be true through the end of the five-year survey, where only HD 186427 b ( d) will end up with full coverage over multiple orbits. Still, as M. R. Giovinazzi et al. (2025) show, NEID data can be combined with archival data over longer baselines to place stringent constraints on the orbits of long-period planets. Further, Figure 12 illustrates that the survey sensitivity is presently limited by the observing baseline, not by the RV precision or the intrinsic variability of the target stars. As the survey progresses, we expect to be able to robustly detect planets with m s-1 out to periods of several hundred days.
6.2 New Candidate Signals
We report new candidate RV signals alongside known signals in Tables 4 and 3. These new signals are described briefly in Section 5, but we refrain from classifying them as either real planets or false positives and we defer detailed analysis to future studies. However, it is worth noting that unlike the known planet signals we recover, which generally have eccentricities consistent with the literature values, many of the new signals we report have fairly eccentric orbits. This is atypical for the known exoplanet population, and may be an outcome of our use of a flat eccentricity prior, which is known to allow for spurious eccentric orbits when the periastron phase is poorly sampled. Any follow-up investigation of these signals should include an analysis of whether the high eccentricities persist when different priors are imposed.
@@ 6.3 Targets on ExEP List
The 2020 Decadal Survey of Astronomy and Astrophysics (National Academies of Sciences, Engineering, and Medicine, 2021) recommended the development and construction of the Habitable Worlds Observatory (HWO), a flagship space telescope with high contrast direct imaging and spectroscopy capabilities from UV to IR wavelengths. This telescope will carry out a survey of nearby stars to search for and and characterize the atmospheres of at least 25 habitable-zone, terrestrial exoplanets. HWO is presently expected to be able to accomplish this through an uninformed search. However, information gleaned from EPRV surveys such as NETS has the potential to greatly increase the efficiency and scientific return of the mission. As R. Morgan et al. (2021) show, partial prior knowledge of Earth-like planet orbits from RV measurements can be used to improve HWO survey yield (i.e., the number of characterized planets) by up to 50% and accelerate the detection of key atmospheric biomarkers. Further, recent work by S. R. Kane et al. (2024) has highlighted the obstacles posed by as-yet undetected massive planets orbiting in or near the habitable zone. Planets as small as can dynamically preclude stable orbits for terrestrial planets through much of the habitable zones of their host stars, rendering these targets unfit for the HWO survey. But for the majority of high priority HWO target stars identified by E. Mamajek & K. Stapelfeldt (2024), existing RV data are insufficient to identify or rule out such perturbers (K. Laliotis et al., 2023; C. K. Harada et al., 2024).
@@
Independent of these challenges, accurate interpretation of atmospheric spectra collected by HWO or similar missions will require precise knowledge of the exoplanets’ physical properties. In particular, insufficient constraints on the planetary mass can lead to large uncertainties in the inferred atmospheric parameters (N. E. Batalha et al., 2019) and incorrect assessments of habitability and of the presence or absence of biosignatures (M. Damiano et al., 2025). M. Damiano et al. (2025) recommend a mass precision of better than 10% to guarantee reliable results from reflected light spectroscopy of Earth-like planets.
@@
The NETS target list includes 23 high-priority HWO target stars, including many with known planets (HD 115617, HD 143761, HD 190360, HD 86728, HD 95735) and others with promising, low-amplitude candidate RV signals identified in this work and previous works (HD 4628, HD 10780, HD 146233). It is essential that these stars continue to be monitored with NETS and other EPRV surveys to guarantee that we have an accurate census of their planetary systems and that these systems are well characterized in preparation for HWO. At the very least, establishing and maintaining an EPRV baseline now will greatly accelerate mass measurements for Earth-like planets independently detected by HWO. NETS data can also serve an important role in continuing to study the impact of stellar variability on the spectra of these host stars; while the NETS targets are quiet, they are not silent, and activity-driven spectral variations are well known to complicate exoplanet spectroscopy measurements (B. V. Rackham et al., 2018, 2019).